# GlobalTomo: A global dataset for physics-ML seismic wavefield modeling and FWI

### v4

@@ 1 Introduction
Global seismic tomography is a crucial yet intricate field within the earth sciences that facilitates understanding the earth’s internal structure and dynamics. This discipline has wide-ranging applications, from the discovery of natural resources (Virieux and Operto, , 2009) and the assessment of seismic hazards (Zang et al., , 2021) to the exploration of our planet’s evolutionary history (Dziewonski and Romanowicz, , 2015). It utilizes seismic signals–earthquake-induced ground vibrations–recorded at surface stations to invert and map the earth’s interior. The fundamental challenge in this field lies in the integration of forward modeling of seismic wave propagation with FWI, both of which are essential for accurately correlating seismic data with velocity structures (Romanowicz, , 2003; Fichtner, , 2010). However, accurately simulating seismic waves through the earth’s complex structures and inverting this data is challenging due to the heterogeneous nature of terrestrial media and the inherently ill-posed characteristics of FWI (Tromp, , 2020; Lyu et al., , 2021).
@@
Modern numerical simulation methods, such as finite difference methods (Kelly et al., , 1976), finite element methods (Zienkiewicz et al., , 2005; Moczo et al., , 2007), and spectral element methods (Komatitsch and Tromp, 2002a, ; Komatitsch and Tromp, 2002b, ; Krischer et al., , 2021), are fundamental in the forward modeling of seismic waves. Although these methods have proven effective, they impose significant computational demands, particularly when deploying higher-resolution models necessary for detailed geophysical interpretation. High-resolution imaging requires the simulation of high-frequency waves, which in turn necessitates smaller time steps and finer spatial discretization. This is essential to prevent numerical dispersion and maintain the stability of the simulations, as mandated by the Courant-Friedrichs-Lewy condition (De Moura and Kubrusly, , 2013). Consequently, as the desired resolution increases, so does the computational burden, exponentially escalating the required computational resources and time, thereby introducing a substantial bottleneck (Li et al., 2022b, ).
@@
The computational intensity of forward modeling significantly constrains the efficiency of FWI. FWI aims to iteratively refine a model of the earth’s structure by minimizing the discrepancies between observed and simulated seismic data (Plessix, , 2006). The prolonged computation times required for forward modeling restrict the exploration of the model parameter space and limit the number of feasible iterations (Bozdağ et al., , 2016; Lei et al., , 2020). This limitation is particularly problematic when attempting high-resolution inversions over extensive areas. Consequently, this situation highlights the critical need for more efficient methodologies in both seismic forward modeling (Nissen-Meyer et al., , 2014) and inversion strategies (Adourian et al., , 2023).
@@
Recent advances in ML have shown significant promise in mitigating the aforementioned computational challenges in seismic tomography. Techniques such as Physics-Informed Neural Network (PINN) (Cuomo et al., , 2022) and Neural Operator Learning (Lu et al., , 2021; Li et al., 2020a, ; Wang et al., , 2021) efficiently model complex physical systems by approximating solutions to the underlying Partial Differential Equations. These methods have been successfully applied in fields including fluid dynamics prediction (Cai et al., , 2021), new material discovery (Vasudevan et al., , 2021), climate change modeling (O’Gorman and Dwyer, , 2018; Pathak et al., , 2022), and enhancing biomedical treatments (Liu et al., , 2020). By significantly reducing computational time compared to traditional numerical simulations, these ML techniques enable more efficient high-resolution seismic modeling and inversion. Furthermore, modern ML approaches can generalize beyond trained scenarios, facilitating the exploration of previously uncharted earth’s interior.
@@
However, the efficacy of ML methods in geophysical applications is heavily contingent upon the availability of extensive, high-quality training data. Currently, there is a significant lack of benchmark synthetic datasets specifically designed for global seismic tomography and the ML community. Although datasets to support FWI (Deng et al., , 2022; Feng et al., , 2023) have been introduced, they are mostly limited to subsurface exploration scenarios and do not fully represent the earth’s entire scale, which is crucial for high-resolution wave propagation modeling. This deficiency highlights a critical gap in the resources needed to advance ML applications in this area.
@@
GlobalTomo’s design is grounded in well-established geological studies (Dziewonski and Woodhouse, , 1987; French and Romanowicz, , 2014; Tromp, , 2020) that provide a credible distribution of earth’s internal properties, ensuring our model is adhering to recognized physical and structural characteristics. For the specific application of earth tomography, we avoid out-of-distribution scenarios, as they involve hypothetical or artificially generated distributions that may lack geological relevance. We employ spherical harmonics to effectively parameterize velocity structures and account for variable seismic sources. Extensive benchmark analyses demonstrate GlobalTomo’s capabilities in fostering generalization, accelerating forward modeling, and addressing the inherent ill-posedness of seismic inversion tasks. By providing GlobalTomo, we aim to bridge the existing gap between geophysical sciences and the ML community, thereby enabling more precise and efficient seismic modeling and inversion processes through ML-driven innovations (Mousavi and Beroza, , 2022).
@@ 1.1 Related Work
FWI has been extensively applied across a variety of domains, including shallow crustal imaging (Tape et al., , 2009; Górszczyk et al., , 2017), continental tomography (Chen et al., , 2007; Zhu et al., , 2012), medical imaging (Schreiman et al., , 1984; Wiskin et al., , 2019), and nondestructive testing (Rao et al., , 2016). In seismology, classical FWI employs numerical simulations to derive seismograms from an initial earth model and iteratively optimizes this model using the adjoint-state method (Tarantola, , 1984; Tromp et al., , 2005). While it has achieved significant success in both local and global inversions (Fichtner et al., , 2009; French and Romanowicz, , 2015), FWI still confronts substantial challenges such as susceptibility to local minima (Tromp, , 2020) and high computational demands especially when Bayesian methods are applied (Yang et al., , 2021). Emerging ML approaches, bolstered by comprehensive global seismic datasets, present promising avenues to address these challenges, potentially leading to more efficient and precise FWI solutions.
@@
Neural operator learning has garnered significant attention for solving parametric PDEs, as evidenced by a range of pioneering studies (Kovachki et al., , 2023; Li et al., 2020b, ; Li et al., 2020a, ; Wang et al., , 2021; Lu et al., , 2021). Among these, DeepONet stands out for its flexibility in modeling mesh-free spaces and producing continuous solutions (Lu et al., , 2021; Wang et al., , 2021). Conversely, the Fourier Operator Learning (FNO) operates with mesh-based inputs and outputs yet efficiently models the frequency domain and demonstrates rapid convergence during training (Li et al., 2020a, ; Lu et al., , 2022). Once trained, neural operators can swiftly predict solutions for new parameter sets with just a single forward pass. This capability is particularly advantageous in seismic applications, enabling the instant modeling of wave propagation under various velocity structures during inference. Such efficiency significantly streamlines the inversion process when compared to PINN methods (Song and Alkhalifah, , 2021; Rasht-Behesht et al., , 2022).
@@ 2.2 Model Configuration
For generating global wavefield and seismogram data for our dataset, we employ AxiSEM3D (Leng et al., , 2019), a forward simulator renowned for its efficiency in simulating global seismic wave propagation. AxiSEM3D is exceptionally adaptable, allowing customization to various model complexities (Leng et al., , 2016). This simulator effectively utilizes the axisymmetry of global wavefields in a source-centered frame and extends calculations into the azimuthal Fourier domain. Further details are discussed in Section B.2.
@@
Each model’s 3D structure within these tiers is developed by introducing perturbations ranging from -10% to 10% to a 1D background model. These perturbations are meticulously designed to mirror the subtle yet significant 3D heterogeneity of the earth’s interior, as uncovered by global tomographic studies (Dziewonski and Woodhouse, , 1987; French and Romanowicz, , 2014; Thrastarson et al., , 2024); even earth’s most extreme features, such as subduction zones and mantle plumes, show velocity variations of less than 10% (Ritsema and Lekić, , 2020; Thrastarson et al., , 2024). Such adjustments are crucial for demonstrating the impact of tomographic features on earth’s geochemical and geodynamical processes, including mantle convection (van der Hilst et al., , 1997), subduction of tectonic plates (Sun et al., , 2014), and deep mantle heterogeneity such as the large low-shear-velocity provinces (Cottaar and Lekic, , 2016). Further details on the configurations of each model are provided in Appendix C.
@@
Perturbations within each defined layer are parameterized using real spherical harmonics (Dahlen and Tromp, , 2020) up to degree 8, with radial values interpolated linearly between layers. Spherical harmonics, employed as the basis for parameterization, are mathematical functions that define a set of orthogonal functions on the sphere. These functions are particularly advantageous in geophysical applications because they can naturally represent the spherical geometry of the earth. Although some current mantle tomographic models offer higher resolution (Ritsema et al., , 2011), spectrum analysis indicates that the significant power predominantly resides at lower degrees (Meschede and Romanowicz, , 2015; Ritsema and Lekić, , 2020). Consequently, our choice to limit structural parameterization to degree 8 in the published dataset effectively captures the predominant long-wavelength heterogeneity of the earth’s interior (Su and Dziewonski, , 1991). This approach reduces the number of parameters and enhances the solvability of this inverse problem.
@@ 2.3 Data Generation
Table 1 details simulation inputs and outputs. Using Latin Hypercube Sampling (LHS), we generated a set of random velocity structures and source configurations for extensive parameter coverage in the area of interest. Simulations were performed on a Dell PowerEdge R740 cluster with Intel Xeon Scalable processors (2,336 cores, up to 2.9GHz, 13,056GB RAM). Forward modeling for Acoustic and Elastic tiers required about 2,800 and 8,400 CPU hours respectively, while seismic simulations for the Real Earth tier needed about 100,000 CPU hours. Despite the high computational demands, this method is more cost-effective than traditional ones (Thrastarson et al., , 2024).
@@
Our dataset comprises two types of output data. The first type is surface seismogram data, which includes one-dimensional time series recordings of ground displacement at each surface station. These seismograms are observable globally and are crucial for the inversion of the earth’s internal structure. The second type is seismic wavefield data, capturing the propagating seismic wavefield within the earth’s interior. Although this spatially dense wavefield cannot be directly measured and used for traditional inversion, it reveals complex interactions between seismic waves and the earth’s three-dimensional structure. The intricacies of the wavefield, which contain rich details, have recently gained attention in global wavefield studies (Leng et al., , 2020) but are often too challenging to identify manually. We provide this intermediate wavefield data to facilitate the training of neural networks. For detailed configurations and distribution of each type of output data, please refer to Appendix C. We compare the related dataset with GlobalTomo from various aspects in Appendix E.
@@ 3.1 Baselines
We implement our baselines using NVIDIA Modulus Sym (Hennigh et al., , 2021), a framework designed for high-performance scientific computing.
@@
Considering the tendency of neural networks to favor low-frequency solutions (Rahaman et al., , 2019), we explore the use of a H-Fourier. This model combines the gating mechanisms of Highway Networks (Srivastava et al., , 2015) with the frequency domain learning capabilities of Fourier Neural Networks (Tancik et al., , 2020). The H-Fourier is designed to handle deep learning tasks requiring both complex layer-to-layer transformations and frequency-specific learning, potentially more adept at managing intricate patterns in seismic data.
@@
Instead of directly mapping velocity structures to fixed output grids, this approach involves building a framework that models both wave propagation and seismogram generation on a continuous spatial-temporal domain. We utilize the DeepONet (Lu et al., , 2021), where the trunk net processes inputs of positions and times, and the branch net handles velocity structures. The outputs from both nets are combined through a dot product before passing through a linear output layer to produce the final prediction. To improve training efficiency, we developed an optimized method for parallel multiplication of the trunk and branch net outputs; see Section F.2.
@@
While DeepONet is trained using standard supervised learning, which might limit its generalization in high-resolution domains, we explore additional capabilities with PIDO (Wang et al., , 2021). This model enforces physically consistent representations by incorporating interior and free-surface constraints during training. These constraints are integrated by calculating the derivatives within the PDEs using automatic differentiation.
@@
The GNOT (Hao et al., , 2023) leverages the strengths of both neural operators and transformer architectures to effectively capture intricate spatial and temporal dependencies in data. By utilizing a self-attention mechanism, GNOT can efficiently model long-range interactions and handle high-dimensional inputs, enabling accurate predictions of dynamic systems like seismic wave propagation.
@@ 3.2.2 Inversion
Figure 5: Inversion with increasing starting points. Correlation between the inverted models and the ground truth strengthens as the number of starting points increases.
@@
An alternative inversion strategy is to directly map observed seismograms to velocity structures. This method bypasses the conventional requirement for a forward ML model, potentially addressing the optimization challenges inherent in traditional FWI. We evaluated this approach by training the InversionNet-3D (Deng et al., , 2022) with a 3D CNN backbone and an InversionMLP with 6 fully connected layers and 1000 hidden units, designed to input acoustic wave seismogram data and output velocity structures. Post-training, these models predicted reasonable velocity structures across various degrees in an unseen test set.
@@ 4 Conclusion
The authors would like to thank Miss Chen Zhen (BIGAI) for making the nice figures, Prof. Yizhou Wang (Peking University) for supporting the project, and NVIDIA for their valuable technical support for accelerated computing and artificial intelligence. S. Li, Z. Mu, S. Xin, and Y. Zhu are supported in part by the National Science and Technology Major Project (2022ZD0114900), the National Natural Science Foundation of China (62376009), the Beijing Nova Program, the State Key Lab of General AI at Peking University, the PKU-BingJi Joint Laboratory for Artificial Intelligence, and the National Comprehensive Experimental Base for Governance of Intelligent Society, Wuhan East Lake High-Tech Development Zone. Z. Li and X. Song are supported in part by the National Key R&D Program of China (2022YFF0800601), China Postdoctoral Science Foundation (2023M740043), and the Emerging Engineering Interdisciplinary–Young Scholars Project, Fundamental Research Funds for the Central Universities. K. Leng is supported in part by the EPSRC grant, Blueprinting for AI for Science at Exascale (BASE-II, EP/X019918/1).

### v5

@@ 1 Introduction
Global seismic tomography is a crucial yet intricate field within the earth sciences that facilitates understanding the earth’s internal structure and dynamics. This discipline has wide-ranging applications, from the discovery of natural resources [69] and the assessment of seismic hazards [75] to the exploration of our planet’s evolutionary history [11]. It utilizes seismic signals–earthquake-induced ground vibrations–recorded at surface stations to invert and map the earth’s interior. The fundamental challenge in this field lies in the integration of forward modeling of seismic wave propagation with FWI, both of which are essential for accurately correlating seismic data with velocity structures [55, 16]. However, accurately simulating seismic waves through the earth’s complex structures and inverting this data is challenging due to the heterogeneous nature of terrestrial media and the inherently ill-posed characteristics of FWI [65, 41].
@@
Modern numerical simulation methods, such as finite difference methods [23], finite element methods [78, 43], and spectral element methods [24, 25, 27], are fundamental in the forward modeling of seismic waves. Although these methods have proven effective, they impose significant computational demands, particularly when deploying higher-resolution models necessary for detailed geophysical interpretation. High-resolution imaging requires the simulation of high-frequency waves, which in turn necessitates smaller time steps and finer spatial discretization. This is essential to prevent numerical dispersion and maintain the stability of the simulations, as mandated by the Courant-Friedrichs-Lewy condition [9]. Consequently, as the desired resolution increases, so does the computational burden, exponentially escalating the required computational resources and time, thereby introducing a substantial bottleneck [37].
@@
The computational intensity of forward modeling significantly constrains the efficiency of FWI. FWI aims to iteratively refine a model of the earth’s structure by minimizing the discrepancies between observed and simulated seismic data [48]. The prolonged computation times required for forward modeling restrict the exploration of the model parameter space and limit the number of feasible iterations [3, 29]. This limitation is particularly problematic when attempting high-resolution inversions over extensive areas. Consequently, this situation highlights the critical need for more efficient methodologies in both seismic forward modeling [45] and inversion strategies [1].
@@
Recent advances in ML have shown significant promise in mitigating the aforementioned computational challenges in seismic tomography. Techniques such as Physics-Informed Neural Network (PINN) [7] and Neural Operator Learning [39, 35, 70] efficiently model complex physical systems by approximating solutions to the underlying Partial Differential Equations. These methods have been successfully applied in fields including fluid dynamics prediction [4], new material discovery [68], climate change modeling [46, 47], and enhancing biomedical treatments [38]. By significantly reducing computational time compared to traditional numerical simulations, these ML techniques enable more efficient high-resolution seismic modeling and inversion. Furthermore, modern ML approaches can generalize beyond trained scenarios, facilitating the exploration of previously uncharted earth’s interior.
@@
However, the efficacy of ML methods in geophysical applications is heavily contingent upon the availability of extensive, high-quality training data. Currently, there is a significant lack of benchmark synthetic datasets specifically designed for global seismic tomography and the ML community. Although datasets to support FWI [10, 15] have been introduced, they are mostly limited to subsurface exploration scenarios and do not fully represent the earth’s entire scale, which is crucial for high-resolution wave propagation modeling. This deficiency highlights a critical gap in the resources needed to advance ML applications in this area.
@@
GlobalTomo’s design is grounded in well-established geological studies [13, 18, 65] that provide a credible distribution of earth’s internal properties, ensuring our model is adhering to recognized physical and structural characteristics. For the specific application of earth tomography, we avoid out-of-distribution scenarios, as they involve hypothetical or artificially generated distributions that may lack geological relevance. We employ spherical harmonics to effectively parameterize velocity structures and account for variable seismic sources. Extensive benchmark analyses demonstrate GlobalTomo’s capabilities in fostering generalization, accelerating forward modeling, and addressing the inherent ill-posedness of seismic inversion tasks. By providing GlobalTomo, we aim to bridge the existing gap between geophysical sciences and the ML community, thereby enabling more precise and efficient seismic modeling and inversion processes through ML-driven innovations [44].
@@ 1.1 Related Work
FWI has been extensively applied across a variety of domains, including shallow crustal imaging [62, 20], continental tomography [5, 77], medical imaging [56, 71], and nondestructive testing [50]. In seismology, classical FWI employs numerical simulations to derive seismograms from an initial earth model and iteratively optimizes this model using the adjoint-state method [63, 66]. While it has achieved significant success in both local and global inversions [17, 19], FWI still confronts substantial challenges such as susceptibility to local minima [65] and high computational demands especially when Bayesian methods are applied [73]. Emerging ML approaches, bolstered by comprehensive global seismic datasets, present promising avenues to address these challenges, potentially leading to more efficient and precise FWI solutions.
@@
Neural operator learning has garnered significant attention for solving parametric PDEs, as evidenced by a range of pioneering studies [26, 36, 35, 70, 39]. Among these, DeepONet stands out for its flexibility in modeling mesh-free spaces and producing continuous solutions [39, 70]. Conversely, the Fourier Operator Learning (FNO) operates with mesh-based inputs and outputs yet efficiently models the frequency domain and demonstrates rapid convergence during training [35, 40]. Once trained, neural operators can swiftly predict solutions for new parameter sets with just a single forward pass. This capability is particularly advantageous in seismic applications, enabling the instant modeling of wave propagation under various velocity structures during inference. Such efficiency significantly streamlines the inversion process when compared to PINN methods [57, 51].
@@ 2.2 Model Configuration
For generating global wavefield and seismogram data for our dataset, we employ AxiSEM3D [32], a forward simulator renowned for its efficiency in simulating global seismic wave propagation. AxiSEM3D is exceptionally adaptable, allowing customization to various model complexities [31]. This simulator effectively utilizes the axisymmetry of global wavefields in a source-centered frame and extends calculations into the azimuthal Fourier domain. Further details are discussed in Section B.2.
@@
Each model’s 3D structure within these tiers is developed by introducing perturbations ranging from -10% to 10% to a 1D background model. These perturbations are meticulously designed to mirror the subtle yet significant 3D heterogeneity of the earth’s interior, as uncovered by global tomographic studies [13, 18, 64]; even earth’s most extreme features, such as subduction zones and mantle plumes, show velocity variations of less than 10% [54, 64]. Such adjustments are crucial for demonstrating the impact of tomographic features on earth’s geochemical and geodynamical processes, including mantle convection [67], subduction of tectonic plates [60], and deep mantle heterogeneity such as the large low-shear-velocity provinces [6]. Further details on the configurations of each model are provided in Appendix C.
@@
Perturbations within each defined layer are parameterized using real spherical harmonics [8] up to degree 8, with radial values interpolated linearly between layers. Spherical harmonics, employed as the basis for parameterization, are mathematical functions that define a set of orthogonal functions on the sphere. These functions are particularly advantageous in geophysical applications because they can naturally represent the spherical geometry of the earth. Although some current mantle tomographic models offer higher resolution [53], spectrum analysis indicates that the significant power predominantly resides at lower degrees [42, 54]. Consequently, our choice to limit structural parameterization to degree 8 in the published dataset effectively captures the predominant long-wavelength heterogeneity of the earth’s interior [59]. This approach reduces the number of parameters and enhances the solvability of this inverse problem.
@@ 2.3 Data Generation
Table 1 details simulation inputs and outputs. Using Latin Hypercube Sampling (LHS), we generated a set of random velocity structures and source configurations for extensive parameter coverage in the area of interest. Simulations were performed on a Dell PowerEdge R740 cluster with Intel Xeon Scalable processors (2,336 cores, up to 2.9GHz, 13,056GB RAM). Forward modeling for Acoustic and Elastic tiers required about 2,800 and 8,400 CPU hours respectively, while seismic simulations for the Real Earth tier needed about 100,000 CPU hours. Despite the high computational demands, this method is more cost-effective than traditional ones [64].
@@
Our dataset comprises two types of output data. The first type is surface seismogram data, which includes one-dimensional time series recordings of ground displacement at each surface station. These seismograms are observable globally and are crucial for the inversion of the earth’s internal structure. The second type is seismic wavefield data, capturing the propagating seismic wavefield within the earth’s interior. Although this spatially dense wavefield cannot be directly measured and used for traditional inversion, it reveals complex interactions between seismic waves and the earth’s three-dimensional structure. The intricacies of the wavefield, which contain rich details, have recently gained attention in global wavefield studies [30] but are often too challenging to identify manually. We provide this intermediate wavefield data to facilitate the training of neural networks. For detailed configurations and distribution of each type of output data, please refer to Appendix C. We compare the related dataset with GlobalTomo from various aspects in Appendix E.
@@ 3.1 Baselines
We implement our baselines using NVIDIA Modulus Sym [22], a framework designed for high-performance scientific computing.
@@
Considering the tendency of neural networks to favor low-frequency solutions [49], we explore the use of a H-Fourier. This model combines the gating mechanisms of Highway Networks [58] with the frequency domain learning capabilities of Fourier Neural Networks [61]. The H-Fourier is designed to handle deep learning tasks requiring both complex layer-to-layer transformations and frequency-specific learning, potentially more adept at managing intricate patterns in seismic data.
@@
Instead of directly mapping velocity structures to fixed output grids, this approach involves building a framework that models both wave propagation and seismogram generation on a continuous spatial-temporal domain. We utilize the DeepONet [39], where the trunk net processes inputs of positions and times, and the branch net handles velocity structures. The outputs from both nets are combined through a dot product before passing through a linear output layer to produce the final prediction. To improve training efficiency, we developed an optimized method for parallel multiplication of the trunk and branch net outputs; see Section F.2.
@@
While DeepONet is trained using standard supervised learning, which might limit its generalization in high-resolution domains, we explore additional capabilities with PIDO [70]. This model enforces physically consistent representations by incorporating interior and free-surface constraints during training. These constraints are integrated by calculating the derivatives within the PDEs using automatic differentiation.
@@
The GNOT [21] leverages the strengths of both neural operators and transformer architectures to effectively capture intricate spatial and temporal dependencies in data. By utilizing a self-attention mechanism, GNOT can efficiently model long-range interactions and handle high-dimensional inputs, enabling accurate predictions of dynamic systems like seismic wave propagation.
@@ 3.2.2 Inversion
Figure 5: Inversion with increasing starting points. Correlation between the inverted models and the ground truth strengthens as the number of starting points increases. Performance on different degrees shows a consistent trend.
@@
An alternative inversion strategy is to directly map observed seismograms to velocity structures. This method bypasses the conventional requirement for a forward ML model, potentially addressing the optimization challenges inherent in traditional FWI. We evaluated this approach by training the InversionNet-3D [10] with a 3D CNN backbone and an InversionMLP with 6 fully connected layers and 1000 hidden units, designed to input acoustic wave seismogram data and output velocity structures. Post-training, these models predicted reasonable velocity structures across various degrees in an unseen test set.
@@ 4 Conclusion
The authors would like to thank Miss Chen Zhen (BIGAI) for making the nice figures, Prof. Yizhou Wang (Peking University) for supporting the project, and NVIDIA for their valuable technical support for accelerated computing and artificial intelligence. Z. Li and X. Song are supported in part by the National Natural Science Foundation of China (42394111) and the China National Science and Technology Major Project (2024ZD1001101). S. Li, Z. Mu, S. Xin, and Y. Zhu are supported in part by the National Natural Science Foundation of China (62376009), the State Key Lab of General AI at Peking University, the PKU-BingJi Joint Laboratory for Artificial Intelligence, and the National Comprehensive Experimental Base for Governance of Intelligent Society, Wuhan East Lake High-Tech Development Zone. Z. Li is supported in part by the China Postdoctoral Science Foundation (2025T180102) and the Emerging Engineering Interdisciplinary Young Scholars Project. K. Leng is supported in part by the EPSRC grant, Blueprinting for AI for Science at Exascale (BASE-II, EP/X019918/1).